Fermi National Laboratory
 

Accelerator Update for January 29, 2007 - January 31, 2007

 

Monday January 29

The day shift began with Operations monitoring the Tevatron (TeV) store 5199 (56.92E30) and the Recycler stash (220.0E10), with the Antiproton Source (Pbar) stacking (18.6E10), with the Linac, Booster, Main Injector (MI), and Recycler operating normally, with MiniBooNE and NuMI off, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

Operations prepared NuMI for access at 9:06 AM.

At 9:52 AM, two Linac RF stations (LRF1 & 2) tripped off. An expert reported that the lower Linac gallery was down. The Duty Electrician reported that a substation 480V breaker had tripped.

A MI expert began a study at 11:29 AM.

Operations established beam to Switchyard (SY) at 11:35 AM for Meson MTest beam commissioning.

The Neutron Therapy Facility (NTF) reported at 1:34 PM that they had finished treating their last patient for the day.

Operations reported at 4:07 PM that the Recycler had stashed antiprotons twice during the day shift.

The MI studies finished at 4:22 PM.

SY commissioning for MTest ended at 5:15 PM.

At 7:14 PM, the Tevatron Electron Lens (TEL1) tripped on a vacuum interlock. It reset okay.

At 10:39 PM, the TEL1 tripped off on another vacuum interlock. It reset okay.

Operations reported that the Recycler had stashed three times during the evening shift.

Tuesday January 30

The midnight shift began with Operations monitoring store 5199 (28.32E30) and the stash (394.1E10), and with Pbar stacking (42.7E10).

At 12:23 AM, an Accumulator RF station (ARF1) tripped off. It reset okay.

Operations terminated store 5199 at 1 AM.

At 1:24 AM, Operations reported that they had run a “Collider Sequencer” program that crashed. Operators discovered that a “close injection helix” command was never issued.

Operations put the TeV through a dry squeeze at 1:34 AM.

At 1:41 AM, Operations reported more sequencer problems. Operators rebooted the Central File Serve System (CFSS).

At 1:54 AM, the TeV suffered a quench when the sequencer crashed on a file share error. Operations contacted experts.

A computer expert reported at 2:55 AM that at 1 AM a process began running on all consoles in the Main Control Room (MCR) with a purpose of freeing up memory. Unfortunately, the process used up a lot of memory and caused the sequencer to crash. He stopped the process.

Operations started cold compressor pump downs at TeV sectors A4, C4, and D1.

Operation put the TeV through a dry squeeze at 5:03 AM.

Operations put the TeV into shot setup at 5:26 AM.

Operations turned off the MI for access at 6:59 AM.

Operations established store 5201 at 7:15 AM with an initial average luminosity of 278.08E30. This store set a New Record Luminosity.

Operations prepared Booster for access at 9:03 AM.

A Preaccelerator expert accessed the H- Source dome at 9:50 AM.

At 12:34 PM, a TeV RF station (TRF8) tripped off.

Linac experts began a study period at 2:03 PM.

At 9:43 PM, the TEL1 tripped off on a vacuum interlock. It reset okay.

At 10:54 PM, the Preaccelerator resistor LCW water-skid tripped off due to a loss of control power. Operators attempted to reset the power, but it tripped off again. Operations called in an expert.

Wednesday January 31

The midnight shift began with Operations monitoring store 5201 (58.87E30) and the stack (143.1E10).

A Preaccelerator expert reported at 12:04 AM that the resistor water-skid had a bad pump. Operations called in LCW experts.

The water experts reported at 3:26 AM that they had replaced the resistor water-skid pump.

Linac resumed sending beam at 3:27 AM.

The Plans for Wednesday

The plans for today are to continue the HV101 replacement. The present store will stay in until Thursday morning.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 01/31/2007    email Fermilab